# LONGER: Scaling Up Long Sequence Modeling in Industrial Recommenders

### v1

@@ 4.3.2. Parameters
We evaluate model capacity by scaling the hidden dimension size while fixing the number of layers to 2 and the input sequence length to 2000. As shown in Figure 5(a), AUC increases steadily with parameter count, following a strong power-law trend ($R^{2}=0.987$). These results demonstrate that widening the model is an effective way to enhance performance under fixed architecture, with no sign of saturation in the current parameter range.
@@ 4.3.3. FLOPs
We analyze model performance by varying the number of layers and sequence length while keeping the model dimensionality fixed at 32. As shown in Figure 5(b), AUC increases steadily with FLOPs, following a strong power-law trend ($R^{2}=0.967$). This indicates that increasing computational resources enables the model to process longer or more complex user behavior sequences, capturing higher-order dependencies and improving prediction accuracy, even with a fixed model width.
@@ 4.4. Online A/B Tests
In this section, we present the results of the online A/B tests, which were conducted to evaluate the effectiveness of the proposed model in real-world scenarios within both Douyin Ads and Douyin E-Commerce Platforms, both are very influential commercial platforms and attracting billions of users. The baseline models in these scenarios are already quite strong, making the observed improvements even more significant. The dual-domain testing allowed us to evaluate the model’s generalization ability in both advertising and e-commerce environments, which are critical components of the platform’s ecosystem.
@@
This section presents the results of the A/B test for Douyin Ads, where we evaluate the performance of our model using two key metrics: ADSS (Advertiser Score) and ADVV (Advertiser Value), which are the most important indicators in industrial advertising systems. The test was conducted across three advertisement formats: Live Streaming, Short Video, and Mall. For Live Streaming, the model resulted in a 1.063% improvement in ADSS and a 1.168% improvement in ADVV. In the Short Video format, ADSS is increased by 2.097%, while ADVV showed a 2.151% improvement. In the Mall format, ADSS is improved by 1.816%, and ADVV was saw a 1.407% increase. These results confirm that the model effectively enhances performance across all advertisement formats with consistent improvements.
@@
4.4.2. Douyin E-Commerce Serivce.
For the A/B test on Douyin E-Commerce, we evaluate the effectiveness of different content formats using two key metrics: Order/U (the number of orders per user) and GMV/U (the gross merchandise volume per user). These metrics help us understand the impact of the model not only on total sales volume but also on user-level engagement and value generation. The results show significant improvements in both metrics. For Live Streaming, Order/U increased by 7.9222%, and GMV/U saw an improvement of 6.5404%, indicating that live streaming contents have a strong positive effect on both the number of orders and the value generated per user. In the Short Video content, Order/U is improved by 4.6125%, and GMV/U is increased by 5.2771%, demonstrating the effectiveness of short video content in boosting overall sales per user. These results highlight the substantial impact of both ad formats, with Live Streaming showing notably larger improvements in both Order/U and GMV/U.
@@ 5. CONCLUSIONS
In this paper, we presented LONGER, a Transformer-based framework designed for efficient and scalable modeling of ultra-long user behavior sequences in industrial recommender systems. By introducing a series design with global tokens, token merge with InnerTrans, hybrid causal attention, and system-level optimizations including the GPU-synchronous framework, mixed-precision and recomputation training, and KV cache serving, LONGER enables end-to-end ultra-long sequence modeling under practical industrial constraints. Extensive experiments on industrial billion-scale datasets and online A/B tests across both advertising and e-commerce domains validate its robustness and generalizability at billion-user industrial scale. Notably, LONGER achieves competitive accuracy while significantly reducing computational overhead, making it well-suited for deployment in latency-sensitive production environments. Future work includes investigating more efficient sequence modeling techniques and improving cross-domain behavior modeling in industry.

### v2

@@ 4.3.2. Parameters
We evaluate model capacity by scaling the hidden dimension size while fixing the number of layers to 2 and the input sequence length to 2000. As shown in Figure 5(a), AUC increases steadily with parameter count, following a strong power-law trend ($R^{2}=0.987$). These results demonstrate that increasing model width effectively enhances performance under fixed architecture, with no sign of saturation in the current parameter range.
@@ 4.3.3. FLOPs
We analyze model performance by varying the number of layers and sequence length while keeping the model dimensionality fixed at 32. As shown in Figure 5(b), AUC increases steadily with FLOPs, following a strong power-law trend ($R^{2}=0.967$). This indicates that increasing computational resources enables the model to process longer or more complex user behavior sequences, capturing higher-order dependencies and improving prediction accuracy, even under a fixed model width.
@@ 4.4. Online A/B Tests
In this section, we present the results of the online A/B tests, which were conducted to evaluate the effectiveness of the proposed model in real-world scenarios within both Douyin Ads and Douyin E-Commerce Platforms, both of which are very influential commercial platforms and attract billions of users. The baseline models in these scenarios are already quite strong, making the observed improvements even more significant. The dual-domain testing allowed us to evaluate the model’s generalization ability in both advertising and e-commerce environments, which are critical components of the platform’s ecosystem.
@@
This section presents the results of the A/B test for Douyin Ads, where we evaluate the performance of our model using two key metrics: ADSS (Advertiser Score) and ADVV (Advertiser Value), which are the most important indicators in industrial advertising systems. The test was conducted across three advertisement formats: Live Streaming, Short Video, and Mall. For Live Streaming, the model achieved a 1.063% improvement in ADSS and a 1.168% improvement in ADVV. In the Short Video format, ADSS is increased by 2.097%, while ADVV showed a 2.151% improvement. In the Mall format, ADSS is improved by 1.816%, and ADVV was increased by 1.407% . These results confirm that the model effectively enhances performance across all advertisement formats with consistent improvements.
@@
4.4.2. Douyin E-Commerce Service.
For the A/B test on Douyin E-Commerce, we evaluate the effectiveness of different content formats using two key metrics: Order/U (the number of orders per user) and GMV/U (the gross merchandise volume per user). These metrics help us understand the impact of the model not only on total sales volume but also on user-level engagement and value generation. The results show significant improvements in both metrics. For Live Streaming, Order/U is increased by 7.9222%, and GMV/U is lifted by 6.5404%, indicating that live streaming contents have a strong positive effect on both the number of orders and the value generated per user. In the Short Video content, Order/U is improved by 4.6125%, and GMV/U is increased by 5.2771%, demonstrating the effectiveness of short video content in boosting overall sales per user. These results highlight the substantial impact of both ad formats, with Live Streaming showing notably larger improvements in both Order/U and GMV/U.
@@ 5. Conclusions
In this paper, we presented LONGER, a Transformer-based framework designed for efficient and scalable modeling of ultra-long user behavior sequences in industrial recommender systems. By introducing a series of architectural designs including global tokens, token merge with InnerTrans, hybrid causal attention, and system-level optimizations including the GPU-synchronous framework, mixed-precision and recomputation training, and KV cache serving, LONGER enables end-to-end ultra-long sequence modeling under real-world industrial constraints. Extensive experiments on industrial billion-scale datasets and online A/B tests across both advertising and e-commerce domains validate its robustness and generalizability at billion-user industrial scale. Notably, LONGER achieves competitive accuracy while significantly reducing computational overhead, making it well-suited for deployment in latency-sensitive production environments. Future work includes investigating more efficient sequence modeling techniques and improving cross-domain behavior modeling in industry.